OpenStudy (anonymous):

medal if you could help me :) A = acute angle cos 1/2 A = √x+1/√2x sin A... ?

OpenStudy (anonymous):

To be clear, is it \[\cos(\frac{ x }{ 2 } )=\sqrt{x}\frac{ 1 }{ \sqrt{2x} }\] or \[\cos(\frac{ x }{ 2 } )=\sqrt{x}\frac{ 1 }{ \sqrt{2}x }\]

OpenStudy (anonymous):

I meant for a plus between the two terms.

OpenStudy (anonymous):

@albertedison98 nah i dont get it :/

OpenStudy (anonymous):

Is the last term one over sqrt(2)*x, or one over sqrt(2x)?

OpenStudy (anonymous):

@albertedison98 from this options , whats the correct one ? a. (√X^2-1)/(√x^2) b. (√X^2)/(√x^2+1) c. √x^2-1 d. √x^2+1 e. (√X^2+1)/(√x)

OpenStudy (anonymous):

one over sqrt(2x)